• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name K. Kumar

    Log entry time 16:05:42 on June20,2004

    Entry number 126416

    keyword=Day shift summary

    8:00 Shift change, left dipole still down
    8:15 Bob will try to bring up dipole slowly
    8:28 Run 2308 ends, 30 microA
    8:29 Run 2309 starts, dipole still being ramped up
    9:05 Bob calls tech to help reset dipole power supply
    9:13 Right arm Q1 trips
    9:14 run 2309 stopped, 81K events, 30 microamps
    10:25 Left D1 and right Q1 still down, Mark Stevens being paged
    11:48 Q1 left has tripped, Todd still downstairs troubleshooting
    12:00 Right D1 coolant level dropping fast; will likely trip
    12:45 Mark Stevens finally contacted, will be here in 15 minutes
    13:40 Mark Stevens has an explanation for all our problems
    14:30 Mark Stevens is waiting for cryotech to figure out pressure drop
    16:00 Mark Stevens and Todd have reset all the magnets and they are being ramped up now.

    Only production runs on this shift are 2308 and 2309, with one dipole down.